OpenStudy (anonymous):

Let f(x) be a fifth-degree polynomial, and g(x) also a fifth-degree polynomial. The degree of f(x) + g(x) must also equal 5. I SAY FALSE? because it dosent have to equal 5 right?

OpenStudy (ash2326):

No, it'd be 5 only I'll explain you with an example

OpenStudy (anonymous):

OK.

OpenStudy (ash2326):

Let \[f(x)= x^5+4x+6\] and \[g(x)=4x^5+4x^4+3x^3+1\] so \[f(x)+g(x)=5x^5+4x^4+3x^3+4x+7\] We can see that f(x)+g(x) also has order 5

OpenStudy (anonymous):

 @ParthKohli am i distributing right?